ligo-ex ligo-ds
  Richardson Lab Experimental Log, Page 8 of 11  Not logged in ELOG logo
ID Date Author Type Category Subject
  187   Tue Aug 1 16:22:35 2023 CelesteUpdateTCSGraphical Updates to HOM RH Modification, Tuesday 8/1
Slides 20-33 are newly added (data/graphs from parameter sweep feature-1/8th heater unit length varied from 10% of original to 100% of original).
Attachment 1: HOM_RH_Simulation_Progress_80123.pdf
HOM_RH_Simulation_Progress_80123.pdf
  186   Tue Aug 1 15:44:31 2023 SophiaUpdateScripts/ProgramsUpdate on Laser Noise Data Collection and Analysis: Dynamic CSD and PSD Plots
We have identified two problems facing the intensity noise calculations, which we are working to solve. The first of these problems, is that the red pitaya appears to be taking data approximately 4 times faster than we would expect from the decimation rate based on how quickly it runs. To see if this was affecting our data, we plugged the red pitaya into a function generator. We found that my code is able to identify the driving frequency, but consistently gives predictions that are roughly 7% greater than we expect. When we drive at 100 Hz, the plot says the peak is at 107 Hz, when driven at 400 Hz, we see between 422-427 Hz, etc. While this implies that the red pitaya is not taking data at a faster rate than we would expect (if it did we would anticipate a lower than expected frequency, by a factor of roughly 4), it does imply some intrinsic issue with either the red pitaya or with the code. To fix this, I am going to use my code on artificially generated signals and see if we still observe these issues.
Attachment 1: Screen_Recording_2023-08-01_at_5.28.39_PM.mov
Attachment 2: Screen_Recording_2023-08-01_at_5.29.46_PM.mov
Attachment 3: Screen_Recording_2023-08-01_at_5.31.00_PM.mov
Attachment 4: Screenshot_2023-08-01_at_5.42.53_PM.png
Screenshot_2023-08-01_at_5.42.53_PM.png
Attachment 5: Screenshot_2023-08-01_at_5.44.34_PM.png
Screenshot_2023-08-01_at_5.44.34_PM.png
Attachment 6: Screenshot_2023-08-01_at_5.45.35_PM.png
Screenshot_2023-08-01_at_5.45.35_PM.png
  185   Tue Aug 1 12:48:41 2023 AidenUpdateCleanroomPre Filter Change

Changed the pre filter inside of the 3 stage HEPA system next to the soldering station. It was rather dirty and I have attached images with a clean filter on the left and the used one on the right. I reset the pre filter age on the system. I tried to see if I could tell if the HEPA filter was dirty but I could not see it. I did not reset the age of the HEPA or UV stations. They are currently 378 days old.

Attachment 1: IMG_4564.jpg
IMG_4564.jpg
Attachment 2: IMG_4563.jpg
IMG_4563.jpg
Attachment 3: IMG_4562.jpg
IMG_4562.jpg
  184   Tue Aug 1 12:10:03 2023 AidenUpdateVACVacuum Chamber Baking Update

Dr. Richardson and I took off the insulation on the lid again to add more insulation in-between the heater tape folds in addition to what was added on top of the tape after discovering discoloration. Also cleaned the lid with IPA wipes.

Also took more RGA data, one with the leak closed and one with the leak open. The graphs below this show that the SNR is improving with the ratio now looking to be around 3. The chamber is still too dirty to be commissioned and we will need to be baked again soon.

Plugged the temperature sensor back in after getting another extension cord. Removed the electronics box for the RGA as well as the full range gauge above the RGA line.

Attachment 1: bake2_230801_ArC.png
bake2_230801_ArC.png
Attachment 2: bake2_230801_ArO.png
bake2_230801_ArO.png
Attachment 3: bake2_230801_Closed_Open.png
bake2_230801_Closed_Open.png
  183   Mon Jul 31 19:27:30 2023 SophiaUpdateScripts/ProgramsUpdate on Laser Noise Data Collection and Analysis: Dynamic CSD and PSD Plots
An update on the CSD/PSD code: The code is now taking a dynamic average of the points, and that has significantly reduced the noise that we were reading and the issues I was having with the lines in the plots. The graph for the CSD is still putting up a fight though, and amassing many lines that do not appear when just scatter plotted. I plan to fix that tomorrow, alongside inserting my own generated signal into the data. I have plotted the coherence alongside these graphs (as shown in the first screen recording and in one of the images), and it is significantly less than 1-- starting at or below 10^(-1). As a note, I am unsure of why the raw data looks like this-- the laser goes through periods of extreme noise and I think this data was taken during one of them.
Attachment 1: Screenshot_2023-07-31_at_10.18.44_PM.png
Screenshot_2023-07-31_at_10.18.44_PM.png
Attachment 2: Screen_Recording_2023-07-31_at_10.16.31_PM.mov
Attachment 3: Screen_Recording_2023-07-31_at_10.19.55_PM.mov
Attachment 4: Screenshot_2023-07-31_at_10.21.44_PM.png
Screenshot_2023-07-31_at_10.21.44_PM.png
Attachment 5: Screenshot_2023-07-31_at_10.26.34_PM.png
Screenshot_2023-07-31_at_10.26.34_PM.png
  182   Mon Jul 31 12:39:01 2023 ShaneUpdateElectronicsLED cable assemblies installed
Today I put together the two LED cable assemblies for the CyMAC and mounted them in the anti-aliasing and anti-imaging chassis.
Attachment 1: IMG_8514.jpeg
IMG_8514.jpeg
Attachment 2: IMG_8516.jpeg
IMG_8516.jpeg
  181   Fri Jul 28 18:03:07 2023 PamellaUpdateFLIRFinal plot - FLIR
[Pamella]
    1. I was able to plot the final result with the data to the heater. I attached below the "3Combined_HighTemp_Gaussian_Plot." in this plot we can see better behavior on the Gaussian compared to the plot in ELOG 169, I was using the same data but with a different approach. On the ELOG 169, we have the center point isolated data and this new plot is the temperature more than 70 C isolated because we have a very good heater temperature distinguish do background. For the all data I got I was using a power current of 0.20A. To get the data I waited for 30 minutes until the heater became stable and after that, I started to take snaps, I took more than one snap for each one different 6 positions on the screen, and We can see the positions on ELOG 167.
    2. Also I attached the calibration plot ("calibration_plot") between the measurements with the FLIR camera and thermocouple and we can see looks good if we compare the final plot.
    3. For better analyses I attached a plot of the calibration line on the Gaussian plot.
    4. To do: I will finish the final report.
Attachment 1: 29Combined_HighTemp_Gaussian_Plot.jpg
29Combined_HighTemp_Gaussian_Plot.jpg
Attachment 2: 3Combined_HighTemp_Gaussian_Plot.jpg
3Combined_HighTemp_Gaussian_Plot.jpg
Attachment 3: calibration_plot(2).png
calibration_plot(2).png
  180   Fri Jul 28 17:45:45 2023 CelesteUpdateTCSGraphical Updates to HOM RH Modification, Friday 7/28
Updated 1/8th Heater simulation modification graphs. Slides 7, 10, 16-19 are new additions.
Attachment 1: HOM_RH_Simulation_Progress.pdf
HOM_RH_Simulation_Progress.pdf
  179   Fri Jul 28 13:46:17 2023 AidenUpdateVACVacuum Chamber Baking Update

Took some more data to look at the HC count. Initial data was from leaving the Argon leak closed and then left it open for 15 minutes and then took data for the open scenario. Attached are the graphs for these where attachment 1 is the closed scenario and attachment 2 is the open scenario.

From these graphs it looks like the signal-noise ratio of our calibration line is poor with the ratio between these two being 1.3. This means that the calibrated leak is barely being detected over the background. This should improve with more baking and pumping. I have also now attached overlayed graphs comparing the chamber when the leak is closed versus when it is open. As well as one comparing the chamber right after the bake versus the most recent data both with the leak open.

Attachment 1: bake2_230720_ArC.png
bake2_230720_ArC.png
Attachment 2: bake2_230720_ArO.png
bake2_230720_ArO.png
Attachment 3: bake2_230728_Closed_Open.png
bake2_230728_Closed_Open.png
Attachment 4: bake2_230728_Open_Open.png
bake2_230728_Open_Open.png
  178   Fri Jul 28 13:39:39 2023 shaneUpdateCleanroomfull 5 zone cleanroom particle count
Here's the 5 zone cleanroom measurement from today.
Attachment 1: download.png
download.png
  177   Fri Jul 28 13:07:39 2023 Shane, Aiden UpdateCleanroomcleanroom cleaning and particle count
[Aiden, Shane]

cleaning cleanroom and particle count

  • 11:08 am: started particle count
    • zone 3:
      • 0.3 u: 4198
      • 0.5 u: 1080
      • 1.0 u: 581
    • zone 4:
      • 0.3 u: 1247
      • 0.5 u: 623
      • 1.0 u: 415
  • 11:24 am: break for removal of vac chamber insulation (pictures attached), with counts before removal as seen above, and counts after removal/replacement for zone 3 as follows:
    • zone 3:
      • 0.3 u: 2993
      • 0.5 u: 415
      • 1.0 u: 207
  • 12:02 pm: began surface check and wipedown
  • 12:13 pm: started vacuuming the floor
  • 12:25 pm: finished vacuuming the floor
  • 12:26 pm: started mopping the floor
  • 12:32 pm: finished mopping the floor
  • 12:32 pm: started cleaning the buckets
  • 12:37 pm: started mopping with IPA wipes
  • 12:43 pm: finished mopping with IPA wipes
  • 12:45 pm: changed sticky floor mats
  • 12:47 pm: started particle count
    • zone 3:
      • 0.3 u: 2951
      • 0.5 u: 290
      • 1.0 u: 41
    • zone 4:
      • 0.3 u: 290
      • 0.5 u: 0
      • 1.0 u: 0
Attachment 1: image_50412289.JPG
image_50412289.JPG
Attachment 2: image_50432769.JPG
image_50432769.JPG
Attachment 3: image_67126017.JPG
image_67126017.JPG
Attachment 4: image_67163137.JPG
image_67163137.JPG
Attachment 5: image_67168001.JPG
image_67168001.JPG
Attachment 6: IMG_8427.jpeg
IMG_8427.jpeg
  176   Thu Jul 27 14:46:55 2023 Shane, JonUpdateElectronicsAA chassis DC on/off switch assembled
Made the first DC on/off switch assembly today and installed in anti-aliasing chassis. Second switch assembly to be completed later.
  175   Thu Jul 27 14:11:42 2023 SophiaUpdateScripts/ProgramsUpdate on Laser Noise Data Collection and Analysis: Dynamic CSD and PSD Plots
Here's a screen recording of the dynamic raw data and PSD/CSD plots as they are now. Will begin working on implementing a code to generate the statistical estimator for the CSD uncertainty now. (Will talk to Tyler more about whether I can get the lines to happen or not)
Attachment 1: Screen_Recording_2023-07-27_at_5.08.25_PM.mov
  174   Thu Jul 27 11:28:36 2023 Sophia UpdateScripts/ProgramsUpdate on Laser Noise Data Collection and Analysis: Dynamic CSD and PSD Plots
This week I have successfully been able to implement our dynamic CSD and PSD plotting code, but along the way I encountered a few challenges which I will outline in this elog post. Challenges: 1. Once I was able to plot the data through time, I produced a graph of the raw data (shown below in the top to images) which showed pulsed data instead of a continuous stream. On Tuesday (26th of July) Tyler and I fixed this issue, realizing that calling the starting the trigger in the red pitaya was the issue. When I commented that line of code out, it immediately started plotting continuously. 2. When that issue was solved, We then started having problems with the automatically generated lines in the python program which would double back on each new iteration. I am still dealing with this issue, because the line likes to follow the progression of the data intake instead of through the plot. Currently, I am trying to plot the data as a scatter plot and then add the line as this might make it care only about the data's distribution instead of the order of the points in the list, but I am still unsure. I might try out np.sort (or some equivalent to sort the entire data set before plotting) as well.
Attachment 1: Screenshot_2023-07-26_at_3.02.05_PM.png
Screenshot_2023-07-26_at_3.02.05_PM.png
Attachment 2: Screenshot_2023-07-26_at_1.38.47_PM.png
Screenshot_2023-07-26_at_1.38.47_PM.png
Attachment 3: Screenshot_2023-07-27_at_1.36.24_PM.png
Screenshot_2023-07-27_at_1.36.24_PM.png
Attachment 4: Screenshot_2023-07-27_at_1.33.01_PM.png
Screenshot_2023-07-27_at_1.33.01_PM.png
  173   Thu Jul 27 10:18:43 2023 CelesteUpdateTCSGraphical Updates to HOM RH Modification
Attached are updated graphs of the irradiance distribution from the 1/8th ring heater units (extruded and reduced in length).
Attachment 1: HOM_RH_1_8_Unit_Modification_Graphs.pdf
HOM_RH_1_8_Unit_Modification_Graphs.pdf
  172   Tue Jul 25 15:35:14 2023 ShaneUpdateElectronicsMore chassis moved
Finished basic assembly of binary input and output chassis today, and moved one to top shelf of work bench. The other is in corner of work bench (image attached). They are delicate, so please do not move them or place anything on top of them.
Attachment 1: IMG_8409.HEIC
  171   Mon Jul 24 16:19:29 2023 Shane, JonUpdateElectronicsChassis semi-assembled and moved
Began assembly on AA and AI chassis today, moved them both to top work bench shelf (image attached). They are delicate so please do not disturb them or put anything on top of them.
Attachment 1: IMG_8386.pdf
IMG_8386.pdf
  170   Mon Jul 24 14:55:53 2023 JonUpdateVACRemoved extension power cable
I removed the extension cable that was powering the leftmost heater controller, since there is no bake currently happening. It is currently powering the LED over the electronics bench, and can be removed if needed. I will be ordering more extension cords to permanentize our setups.
  169   Thu Jul 20 21:00:23 2023 PamellaUpdateFLIRData plot FLIR and updates about the black body
[Pamella]
  • Today I got more data to plot the Gaussian. So I took more snaps in each position for the six different spots than we be using to have a better calibration of the FLIR collected data. I attached the new plot below. Also, I did the same plot for each region as on the Elog 167 but I have more than 20 pics because I was using a big number of data so I just attached one example below.
  • To access the Elog, click here.

  • Also I think we have some real (non-ideal) heat diffusion by the screen and not noise like Dr. Richardson suggested. I was testing today and we can see the first pic before the start heater source turns on, the second pic is at 120.2 C (0.31A) with the heater on and the last pic is after the heater source cooled back down to room temperature. Just in the second pic, we have a strong spot on the top, so it looks like a non-ideal diffusion.
Attachment 1: AcquisitionImage(Jul-20-2023_13_09).jpg
AcquisitionImage(Jul-20-2023_13_09).jpg
Attachment 2: AcquisitionImage(Jul-20-2023_13_15).jpg
AcquisitionImage(Jul-20-2023_13_15).jpg
Attachment 3: AcquisitionImage(Jul-20-2023_13_33).jpg
AcquisitionImage(Jul-20-2023_13_33).jpg
Attachment 4: AcquisitionData-480(Jul-20-2023_SIsolated_Plot.png
AcquisitionData-480(Jul-20-2023_SIsolated_Plot.png
Attachment 5: Gauss.plot-2.jpeg
Gauss.plot-2.jpeg
  168   Thu Jul 20 18:03:19 2023 AidenUpdateVACVacuum Chamber Baking Update

Took another set of RGA data after the second bake has cooled down to 27 degC.

Just by visual comparison to the other graphs, the HC levels of the chamber has gone down further. I will soon overlay the data for an easier comparison.

I also reinstalled the full range gauge above the RGA line. It is still adjusting itself downwards so I will have to check again to get an accurate measurement of the pressure.

Attachment 1: bake2_230720.png
bake2_230720.png
  167   Wed Jul 19 19:10:31 2023 PamellaUpdateFLIRData plot FLIR and updates.
[Pamella]
  • Today I was able to plot a graph for the isolation point on the center of the heater. I got data from six different positions on the screen (I shifted the all coordinates). I extracted the data for the center point and plot the Gaussian with this extracted data for temperature. I attached the all plots below
  • Also I took a snap using the black wall and with the heater at 120.1 C (0.30 A) try to have less noise but we can see this is not very good. At this temperature, we have noise on the top and I don't understand why because the heater is not in this location. I attached a snap below.
Attachment 1: 1_Isolated_Plot.png
1_Isolated_Plot.png
Attachment 2: 2_Isolated_Plot.png
2_Isolated_Plot.png
Attachment 3: 3_Isolated_Plot.png
3_Isolated_Plot.png
Attachment 4: 4_Isolated_Plot.png
4_Isolated_Plot.png
Attachment 5: 5_Isolated_Plot.png
5_Isolated_Plot.png
Attachment 6: 6_Isolated_Plot.png
6_Isolated_Plot.png
Attachment 7: Gauss.plot-2.jpeg
Gauss.plot-2.jpeg
Attachment 8: AcquisitionImage(Jul-19-2023_17_01).jpg
AcquisitionImage(Jul-19-2023_17_01).jpg
  166   Wed Jul 19 14:39:19 2023 Pamella, Shane, AidenUpdateCleanroomcleaning and particle count
[Pamella, Shane, Aiden]

cleaning cleanroom and particle count

  • 1:10 pm: started particle count
    • Zone 3:
        0.3 u: 2535
        0.5 u: 207
        1.0 u: 41
    • Zone 4:
        0.3 u: 789
        0.5 u: 41
        1.0 u: 0
  • 1:30 pm: began surface check and wipedown
  • 1:41 pm: started vacuuming the floor
  • 1:55 pm: finished vacuuming the floor
  • 1:57 pm: started mopping the floor
  • 2:01 pm: finished mopping the floor
  • 2:02 pm: started cleaning the buckets
  • 2:08 pm: started mopping with IPA wipes
  • 2:17 pm: finished mopping with IPA wipes
  • 2:18 pm: changed sticky floor mats
  • 2:19 pm: started particle count
    • Zone 3:
        0.3 u: 3990
        0.5 u: 1080
        1.0 u: 374
    • Zone 4:
        0.3 u: 540
        0.5 u: 207
        1.0 u: 0
  165   Tue Jul 18 16:35:06 2023 PamellaUpdateFLIRPlot-Data with the new configuration without the mask and reflector.
[Pamella]
  • I was able to plot the first graph for a result between the six different positions on the screen, for now we can see the behavior of the heater temperature in a Gaussian graph with combination data between the six files.
  • To do: Tyler gave me some ideas today to improve the plot. So I'm going to change the code to have insulation on the values for just the heater ("insulation") and I'm going to plot after this insulation data as well I'm going to get more data and compare with more data for the same position.
  • I was using the data than I got last week and I shared on Elog (151) and we can see on this quote

Quote:
[Pamella] I started collecting data by moving the camera FLIR and covering all six positions on the screen (2x3). I was looking and taking snaps for more than one hour and a half. I'll repeat this one more time to make sure we have enough data to do analysis. I attached snap data below. This images is processed on python code.

 

Attachment 1: Gauss.plot.jpeg
Gauss.plot.jpeg
  164   Tue Jul 18 16:12:27 2023 PamellaUpdateFLIRThermocouple plot.
  • I noticed that these measurements have some issues with the weather on different days. We can see in the photo attached below how different the temperatures are on different days, I took the data with the same procedure every day, but we can see the differences between them.
  • To Do: I will do a new data collection using FLIR and thermocouple at the same time to plot comparison between both.

Quote:
[Pamella]
  • I collected data to plot a calibration with the heater. I took measurements with current and temperature (the thermocouple - thermometer) to compare with the FLIR measurements.
  • I made a plot with this data and we can see how temperature vs current behaves. Note: This data I measured manually.

 

Attachment 1: Comp.plot.jpeg
Comp.plot.jpeg
  163   Tue Jul 18 16:04:26 2023 PamellaUpdateFLIRSetup FLIR with black wall (black body)
[Pamella]
  • Today I tried get data the heater with the black screen but doesn't looks possible have just one "energy" point straight to FLIR camera. Tyler and I tried different current and temperatures but keep very bad data. I attached a snap below.
  • I attached a photo about the new setup below. The FLIR is in the most close point possible/safety with the heater. The heater is very close to the black wall but is not touch the screen so is safety.
Attachment 1: AcquisitionImage(Jul-18-2023_15_24).jpg
AcquisitionImage(Jul-18-2023_15_24).jpg
Attachment 2: IMG_8721.jpeg
IMG_8721.jpeg
Attachment 3: IMG_8722.jpeg
IMG_8722.jpeg
  162   Tue Jul 18 14:13:46 2023 Sophia UpdateScripts/ProgramsUpdate on Laser Noise Data Collection and Analysis
We [Sophia and Tyler] have found a code which will partition our data and generate PSD's and CSD's automatically. We are now in the stage of writing this into a dynamic code, which so far has been fast-paced because of the framework we have built for it over the past few weeks. Something to note is that we have found, especially over long periods of time, the laser goes through periods of extreme noise and the signals often cross from where we set them at some point throughout our data collection. We have made the assumption that this is something to do with the laser itself, and not our setup or surrounding space because of both the times that these noise/intensity spikes took place and the duration of the spikes (1.5-2.5 hours). To try and correct for this, we have been "cutting" the noise, and just looking at sections which don't include these big fluctuations. An example of the "noise cut data" is shown in the graph titled "Signals Over Time 7-18-23 Noise Cut Data" and all of the PSD/CSD graphs have been made using only these "noise cut" regions.
Attachment 1: 7_14_23_cut_analyzed_graph.png
7_14_23_cut_analyzed_graph.png
Attachment 2: 7_14_23_raw_graph.png
7_14_23_raw_graph.png
Attachment 3: 7_18_23_cut_analyzed_graph.png
7_18_23_cut_analyzed_graph.png
Attachment 4: 7_18_23_raw_graph_(1).png
7_18_23_raw_graph_(1).png
Attachment 5: 7_18_23_cut_raw_graph.png
7_18_23_cut_raw_graph.png
Attachment 6: 7_18_23_cut_just_PSD_analyzed_graph.png
7_18_23_cut_just_PSD_analyzed_graph.png
  161   Mon Jul 17 14:01:41 2023 AidenUpdateVACVAC System Heating Update

The insulation on the lid of the vacuum chamber has some discoloration and it may be wise to place a sensor closer to this spot in future bakes.

Attachment 1: IMG_4535.jpg
IMG_4535.jpg
  160   Mon Jul 17 13:50:41 2023 AidenUpdateVACVAC System Heating Update

Started to cool down the chamber today. Final temperature readings of the chamber were 146C and 91C for the body and cross respectively.

  159   Mon Jul 17 11:46:24 2023 PamellaUpdateFLIRThermocouple plot.
[Pamella]
  • I collected data to plot a calibration with the heater. I took measurements with current and temperature (the thermocouple - thermometer) to compare with the FLIR measurements.
  • I made a plot with this data and we can see how temperature vs current behaves. Note: This data I measured manually.
Attachment 1: calibration_plot-2.jpeg
calibration_plot-2.jpeg
Attachment 2: calibration_plot-3.jpeg
calibration_plot-3.jpeg
  158   Sun Jul 16 15:21:56 2023 PamellaUpdateGeneralGeneral update about some problems in the lab.
[Pamella]
    General information about the lab facility.
    1. Changed sticky floor mats, because close to entrance the cleanroom, both sticky floor had a many died ants on top.
    2. The light on top of the internet cable bridge is burned out.. I just saw this today but I am not sure if was like this before.
    3. The ant bait traps seem very efficient and right now only a few ants are running around in the lab, most are dead in the bait traps, so probably in a day or two we can change those bait traps.
Attachment 1: IMG_8637.jpg
IMG_8637.jpg
Attachment 2: IMG_8633.jpg
IMG_8633.jpg
Attachment 3: IMG_8631.jpg
IMG_8631.jpg
  157   Fri Jul 14 16:49:42 2023 PamellaUpdateFLIRNew bridge and calibration data
[Pamella]
    Aiden 3D printed a new bridge for the heater and I installed the new bridge yesterday.
    I started collecting data to plot a calibration with the heater. I'm doing measurements with current and the thermocouple (thermometer) to compare with FLIR measurements and have a good calibration.
Attachment 1: IMG_8606.jpeg
IMG_8606.jpeg
  156   Fri Jul 14 14:37:25 2023 AidenUpdateVACVAC System Heating Update

Checked on the vacuum system and it looks like it will either need to be increased in temperature or baked for a longer time. The main body is at 146C while the cross is at 91C.

  155   Thu Jul 13 16:59:03 2023 JonUpdateGeneralAnt traps placed around lab

Update: After seeing the strong response to the first set, I redistributed them to the hottest spots and added one more trap.

Quote:

In response to reports that ants have been observed in the lab, I placed five ant bait traps around the room today. Each is sitting on the floor on top of a piece of aluminum foil, in areas unlikely to be inadvertently stepped on. They contain liquid which will spill out if picked up, so please take care not to disturb them.

 

  154   Thu Jul 13 15:31:21 2023 JonUpdateGeneralAnt traps placed around lab

In response to reports that ants have been observed in the lab, I placed five ant bait traps around the room today. Each is sitting on the floor on top of a piece of aluminum foil, in areas unlikely to be inadvertently stepped on. They contain liquid which will spill out if picked up, so please take care not to disturb them.

  153   Thu Jul 13 14:56:43 2023 AidenUpdateVACVAC System Heating Update

Checked on the vacuum system today to change the PID controls back to their default values in order to try and get the temperature of the cross in front of the turbo pump higher. Currently it is sitting at 90 degC before changing the controls. While the main body is at 146 degC.

Will update again tomorrow on new equilibrium with the default PID settings.

  152   Mon Jul 10 13:23:23 2023 AidenUpdateVACVAC System Heating Update

Checked on the system today.

The flange on the turbo pump has reached an equilibrium temperature of 87 degC when set at 150 degC. While the main body of the chamber has reached an equilibrium temperature of 142 degC.

  151   Sat Jul 8 14:26:17 2023 PamellaUpdateFLIRData with the new configuration without the mask and reflector.
[Pamella] I started collecting data by moving the camera FLIR and covering all six positions on the screen (2x3). I was looking and taking snaps for more than one hour and a half. I'll repeat this one more time to make sure we have enough data to do analysis. I attached snap data below. This images is processed on python code.
Attachment 1: region2-3.jpg
region2-3.jpg
Attachment 2: region3-2.jpg
region3-2.jpg
Attachment 3: region4.jpg
region4.jpg
Attachment 4: region5.jpg
region5.jpg
Attachment 5: region6.jpg
region6.jpg
Attachment 6: region_temperature1-2.jpg
region_temperature1-2.jpg
  150   Fri Jul 7 15:11:39 2023 AidenUpdateVACVAC System Heating Update

The temperature of the flange connecting to the turbo pump after reaching equilibrium is 74 degC. This means that it is safe to proceed with raising the temperature from 120 degC to 150 degC.

Raised the set temperature of the chamber from 120 degC to 150 degC in 15 deg steps.

  149   Thu Jul 6 18:03:19 2023 PamellaUpdateFLIRFLIR project - New configuration
[Pamella]
    Cao suggested using the project without the reflector and mask because in this case we probably got good images around the screen. So I started this yesterday and today I started to get some data to see the stabilization position and check how the parameters fluctuated. I was looking and taking snaps for an hour and a half. I'll repeat this one more time to make sure we have enough data to do analysis.
    To do: The next step is to start collecting data by moving the camera FLIR and covering all six positions on the screen (2x3). Also Aiden is on hand to help and is going to make a new 3D print bridge to have good heater support.
    I attached a image for the new configuration below and a snap data.
Attachment 1: IMG_8193.jpeg
IMG_8193.jpeg
Attachment 2: IMG_8183.jpeg
IMG_8183.jpeg
Attachment 3: AcquisitionImage(Jul-06-2023_17_18)_(1).jpg
AcquisitionImage(Jul-06-2023_17_18)_(1).jpg
Attachment 4: region_temperature1.jpg
region_temperature1.jpg
  148   Thu Jul 6 16:16:03 2023 AidenSummaryVACVacuum Chamber Baking Update

Pressure Measured Before Removing Gauge 2.31 E-7 Torr.

Went in and turned on the heating tape in increments of 30 degC until 120 degC was reached. I will let it equilibrate over night and then asses whether or not the chamber can be raised to 150 degC without having the flange to the turbo pump reach over 120 degC as it is not recommended for that flange to be any hotter.

If it can safely be raised to 150 degC, then only 2 days of baking is necessary. If 120 degC is the bake temperature, a week will be needed.

  147   Thu Jul 6 13:34:48 2023 Pamella, Shane, Julian, AidenUpdateCleanroomcleaning and particle count
[Pamella, Julian, Aiden, and Shane]

    cleaning cleanroom and particle count

  • 12:00 pm: started particle count
    • Zone 3:
        0.3 u: 2826
        0.5 u: 457
        1.0 u: 166
    • Zone 4:
        0.3 u: 457
        0.5 u: 83
        1.0 u: 41
  • 12:17 pm: began surface check and wipedown. NOTE: found a few ants crawling up the metal table frame and on the desk
  • 12:29 pm: started vacuuming the floor
  • 12:45 pm: Finished vacuuming the floor
  • 12:46 pm: Started mopping the floor
  • 12:52 pm: Finished mopping the floor
  • 12:55 pm: Started cleaning the buckets
  • 1:02 pm: Started mopping with IPA wipes
  • 1:11 pm: Finished mopping with IPA wipes
  • 1:12 pm: Changed sticky floor mats
  • 1:14 pm: started particle count
    • Zone 3:
        0.3 u: 2577
        0.5 u: 457
        1.0 u: 166
    • Zone 4:
        0.3 u: 457
        0.5 u: 41
        1.0 u: 0
  146   Wed Jul 5 16:02:12 2023 AidenUpdateClean & BakeClean and Bake batch 14 Bagging
Bagged and Tagged the 2.75" copper gaskets from Batch 14 and placed them into the clean room on the work bench.
  145   Wed Jul 5 15:46:47 2023 Aiden, CaoInfrastructureVACVacuum Chamber Baking Update

Cao and Aiden put the RGA back on to the chamber and measured the outgassing rate after the chamber has cooled down from its first bake. The figure attached shows the RGA measurements with the Argon leak open.

Also put the Full Range Gauge #2 [RGA line] back onto the system to measure the pressure and got a reading of 2.78 E-7 Torr. We believe it is actually lower as the chamber temp is still at 28 degC rather than the 24 it was previously measured at. The gauge also may have needed more time to equilibrate.

Attachment 1: bake1_230705.png
bake1_230705.png
  144   Tue Jul 4 15:35:38 2023 PamellaUpdateFLIRFLIR - reflexion problem
[Pamella] I was testing all the positions we need to cover when moving the FLIR camera or the reflector. I tried to find out more about the problem so I tested the position without the mask, just with the old bridge and I got the same problems, in the middle of the screen we have a very good image but if we move we have the visualization problem and we don't see the reflection very well (intensity, emissivity) in FLIR. For this test I used very low current (0.04 A) for keep everything safety. I also think that problem is not in the focal point of the reflector because we can see very well in the center of the screen, but maybe it is because in the parabolic reflector the reflection in the focal point is "extremely" straight, so for you to see the reflection in this point you need keep close to the middle of the screen, to see around the reflector, at the extremes, we need to change the position for the cable or the light to "change" how the flashes of light reflect and thus achieve exactly the position on the camera. It is something like the sketch/ photo attached below.
Attachment 1: AcquisitionImage(Jul-03-2023_14_29).jpg
AcquisitionImage(Jul-03-2023_14_29).jpg
Attachment 2: AcquisitionImage(Jul-03-2023_14_31).jpg
AcquisitionImage(Jul-03-2023_14_31).jpg
Attachment 3: AcquisitionImage(Jul-03-2023_14_49).jpg
AcquisitionImage(Jul-03-2023_14_49).jpg
Attachment 4: Caderno_sem_título_(26)-2.jpg
Caderno_sem_título_(26)-2.jpg
Attachment 5: parabol1.print.png
parabol1.print.png
  143   Mon Jul 3 18:38:13 2023 PamellaUpdateFLIRFLIR project
[Pamella]
  • I was able to work on the python code to do analysis on the FLIR-Reflector data.
  • I could plot images of total area (Tyler help me on that) with csv file and also could have isolation area for triangles. I have attached examples below. Just for now the images are of different data, so there are some differences in the shapes.
  • I'm working on getting a complete analysis code to work with the different positions of the triangles and to be able to do out the statistical analysis.
  • Also we keeping have problem to get good data if when we move the camera or the reflector on horizontal or vertical position. I am working a some ideas for that.
Attachment 1: Region.jpg
Region.jpg
Attachment 2: triangular_region_temperature-2.jpg
triangular_region_temperature-2.jpg
Attachment 3: heatmap.jpg
heatmap.jpg
Attachment 4: isolation1-4.jpg
isolation1-4.jpg
  142   Mon Jul 3 16:26:21 2023 CaoInfrastructureDAQAccess and control Red Pitaya remotely and run it on local machine
[Cao]

To access and control the Red Pitaya using Python locally on a machine within the local network, one should follow these steps:

  1. Start the SCPI server. This is achieved by first log onto the Red Piatay page

    rp-xxxxxx.local/ 
  2. Go to Development >> SCPI server and turn the server on. (Note : The server is currently running)
  3. Communication with Red Piataya is done through PyVista, install PyVista with:
     sudo pip3 install pyvisa pyvisa-py 
    This has been installed on Chimay. Ensure that you have pip3 install, if not, you can install it using:
     sudo apt-install python3 pip 
  4. To start talking to the RedPitaya, ensure you have the scipt
     redpitaya_scpi.py 
    in your local folder. This is the standard class that you will import to your code to establish connection with the Red Pitaya. This code can be found in directory
     ~/RedPiatya 
     or this link
    

  141   Mon Jul 3 14:32:24 2023 CaoInfrastructureVACCooling down vacuum chamber
[Cao]

After leaving the the chamber to bake at 120 deg C from Thursday 3:41 pm. Today I started to cool the chamber down to room temperature at 11:30 am (total bake duration: 91 hours 49 minutes). Ideally, this process should be ramped down slowly, approx. 6 degree per hour. However we had no ramping function with out controller. The only method to tune this cool/ heat rate is to tune the PID parameters, which are:

  1. Pb : Proportional band, this quantity is inversely proportional to the P-gain
  2. ti : Integral time. Increasing integral time makes the output response slower to error, thus the opposite effect of increasing integration gain
  3. td : Differential time
Since the differential term doesn't contribute to rise time significantly, we neglect it for now and first change Pb and ti . The factory default settings :
  1. Pb : 50
  2. ti : 100 s
  3. td : 25 s
Initially Ichanged Pb from 50 to 200 (max) and ti to 999 s (max) but there is a large offset error. Currently, Pb and ti are sitting on the following settings:
  1. Pb : 200
  2. ti : 800 s

At this setting it takes 25-30 mins to drop 6 degrees, which is not as slow as we want but we will have to tune our PID more carefully. At the first stage, I dropped the temperature to 60 deg C. After 3 hours, we dropped the setpoint to 25 degrees and let the vacuum chamber to cool to room temperature.
  140   Fri Jun 30 16:43:14 2023 PamellaUpdateFLIRUpdates: Small moving the FLIR camera
[Pamella]
    Today, I was able to make some adjustments to the FLIR camera angle, suggested by Dr. Richardson, and we got small possible movements keeping the same pattern in the shapes of the triangles. I have attached photos below. I was able to move left and right as well as up and down and it worked in four points that I marked on the table.
    The only problem is that it only worked on a part of the screen, if I try to place it farther to the right or to the left, we are back to the same problem as yesterday.
Attachment 1: AcquisitionImage(Jun-30-2023_16_23)_(3).jpg
AcquisitionImage(Jun-30-2023_16_23)_(3).jpg
Attachment 2: AcquisitionImage(Jun-30-2023_16_23)_(1)(1).jpg
AcquisitionImage(Jun-30-2023_16_23)_(1)(1).jpg
Attachment 3: AcquisitionImage(Jun-30-2023_16_24).jpg
AcquisitionImage(Jun-30-2023_16_24).jpg
Attachment 4: AcquisitionImage(Jun-30-2023_16_22)_(1)(1).jpg
AcquisitionImage(Jun-30-2023_16_22)_(1)(1).jpg
Attachment 5: AcquisitionImage(Jun-30-2023_16_22)_(1).jpg
AcquisitionImage(Jun-30-2023_16_22)_(1).jpg
  139   Fri Jun 30 15:57:28 2023 PamellaUpdateFLIRData and python code update
    [Pamella]
    Updates: Problems with the emission intensity and python code.
  • Yesterday I was working on get data from FLIR reflector but unfortunately we got some problems:
    1. I realized than if I move the reflector a to left or right the screen doesn't get data very well(I attached a photo below). This is a problem because our idea is have the same type of emission every part on screen. Tyler and I worked to tried fix that but don't had success.
    2. Dr. Richardson gave me the idea to move the FLIR camera to left or right (The same happens if I move up and down) and keep the refletor on the same position every time but unfortunately we got the same problem, the screen doesn't get data very well(I attached a photo below). Now I am working to tried fix that.
  • Also I was able to work on the code to isolate the triangle shape for analyzes, I attached the image for that below.
Attachment 1: isolation1-2.jpg
isolation1-2.jpg
Attachment 2: AcquisitionImage(Jun-29-2023_18_27)_(1).jpg
AcquisitionImage(Jun-29-2023_18_27)_(1).jpg
Attachment 3: AcquisitionImage(Jun-29-2023_18_38).jpg
AcquisitionImage(Jun-29-2023_18_38).jpg
Attachment 4: AcquisitionImage(Jun-29-2023_18_39).jpg
AcquisitionImage(Jun-29-2023_18_39).jpg
  138   Thu Jun 29 18:27:26 2023 Jon, CaoInfrastructureVACVacuum bake attempt 3: success!
[Jon, Cao]

Summary : We resolved problems with heaters tripping power and were able to proceed with chamber baking

1. Circuit connection adjustment

After yesterday elog 137, today we resolved most of these issues. After Jon contacted Facilities, the LP3B 6 circuit was reset and the cleanroom filter & light panels resumed to work as normal.

Regarding the connection of the heaters. we made the following adjustments:

  • High-temperature controller powering lid + upper volume heater: connect to LP3B 8 circuit (clean room sides, 2 outlets)
  • High-temperature controller powering bottom + lower volume heater : connect to LP3B 4 circuit (workstation side, 2 outlets)

2. Replacement of vacuum nipple insulation

We had also received new insulation pieces from Worbo today to replace the existing insulation with the new ones (see images). These cover the 2 4" tubes (for 6" flanges ) and the 4 1.5" tubes (for 2.75" flanges). The new insulations fit perfectly on these tubes. I also placed all insulation taken off from the last elog back onto the chamber ( these are insulations for the pumps and RGA lines).

3. Baking

We started ramping up chamber temperature at 1:41 pm over the course of two hours:
  • Starting set point: 40 deg C
  • Step increase: 10 deg C up to 80 deg C, 5 deg C from 80 deg to 120 deg C
  • At each step, the temperature readouts show approx. 2.1 deg C overshooting, wait to settle back to approx 1.5 deg C overshoot before increase the set point again
We noticed some smoke emanating from flexible bypass line insulation but none from other locations that were previously smoking . We think this is because the high winding density around the bypass line for such a small volume. Adjustment for next bake: Change the location of PID controller RTD to the bypass line . For now, we remove the insulation around bypass to prevent insulation overheating and encourage convection cooling (see image)

The temperature of the chamber reached as stable 120 deg C without any power issues at 3:45 pm. I waited another 15 minutes to verify its stability and the official baking duration started at 4:00 pm Jun 29 2023. Since we are baking at 120 deg C instead of the standard 150 deg C for Aluminium, the duration for the bake will be over three days until Monday morning, upon which we will slowly ramp down the the temperature.

Attachment 1: NewInsulation.png
NewInsulation.png
Attachment 2: BypassLineInsulation.jpg
BypassLineInsulation.jpg
ELOG V3.1.3-7933898